Rowena's age is three more than twice Mary's age. three years ago, Rowena was four times as old as mary was then. How old are they now?
Answer by ankor@dixie-net.com(22740)

Rowena's age is three more than twice Mary's age.
three years ago, Rowena was four times as old as mary was then.
How old are they now?
:
Let r = R's age
let m = M's age
:
WRite an equation for each statement.
:
"Rowena's age is three more than twice Mary's age."
r = 2m + 3
" three years ago, Rowena was four times as old as mary was then."
r - 3 = 4(m-3)
r - 3 = 4m - 12
r = 4m - 12 + 3
r = 4m - 9
Replace r with (2m+3), from the 1st statement
2m + 3 = 4m - 9
3 + 9 = 4m - 2m
12 = 2m
m = 12/2
m = 6
Mary is 6 yrs old
then
r = 2m + 3
r = 2(6) + 3
r = 15 yrs is Rowena's age
:
;
See if that checks out in the statement:
" three years ago, Rowena was four times as old as mary was then. "
15 - 3 = 4(6 - 3)
12 = 4(3)